Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

MGMAT CAT Problem

Expert replies
by thp510 » Sun Mar 06, 2011 2:12 pm
Can anyone show me how to solve this problem WITHOUT using combinatorics and instead, just using probability/fractions?

A small company employs 3 men and 5 women. If a team of 4 employees is to be randomly selected to organize the company retreat, what is the probability that the team will have exactly 2 women?

OA: [spoiler]3/7[/spoiler]
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Sun Mar 06, 2011 3:23 pm
thp510 wrote:Can anyone show me how to solve this problem WITHOUT using combinatorics and instead, just using probability/fractions?

A small company employs 3 men and 5 women. If a team of 4 employees is to be randomly selected to organize the company retreat, what is the probability that the team will have exactly 2 women?

OA: [spoiler]3/7[/spoiler]
The probability that an event will occur exactly n times = P(one way) * total number of possible ways.

One way to choose exactly 2 women is to select first 2 women and then 2 men:
P(WWMM) = 5/8 * 4/7 * 3/6 * 2/5 = 1/14.

The result above needs to be multiplied by the total number of ways to choose exactly 2 women:
WWMM
WMWM
WMMW
MWWM
MWMW
MMWW

Since there are 6 ways to choose exactly 2 women:
6 * 1/14 = 3/7.

We can determine the answer more efficiently if we use combinatorics to count the number of ways 2 women can be chosen. The list above represents the total number of ways to arrange the letters WWMM: 4!/2!2! = 6.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by anshumishra » Sun Mar 06, 2011 3:50 pm
Using combinatorics :
probability = # of ways to select 2 women and 2 men / # total no. of ways to select 4 persons = (3C2*5C2) / 8C4 = 3*10/70 = 3/7
Thanks
Anshu

(Every mistake is a lesson learned )
Join the discussion

by thp510 » Mon Mar 07, 2011 8:07 am
Great explanations. Thanks guys
Join the discussion